# End-to-End Probabilistic Framework for Learning with Hard Constraints

### v1

@@ 5 Conclusion
In this work, we propose ProbHardE2E, which seamlessly incorporates constraints ranging from commonly used linear constraints to challenging nonlinear constraints into black-box probabilistic deep learning models using a differentiable probabilistic projection layer (DPPL). We show that ProbHardE2E is applicable in a wide range of scientific and operational domains, ranging from linear coherency constraints in time series forecasting to nonlinear conservation laws in solving PDEs. Contrary to prior works (Hansen et al., 2023), which only support linear global conservation constraints, ProbHardE2E can support nonlinear conservation constraints. This paves the way for future work to enforce conservation locally over sub-regions or over control volumes à la finite volume methods (LeVeque, 2002). Future work also includes extending our method to handle general non-convex, nonlinear inequality constraints using advanced optimization techniques or relaxations, and to richer covariance parameterizations, e.g., low-rank or dense.
@@ Appendix C Computation of Posterior Distribution for Various Constraint Types
Table 5: Summary of DPPL in ProbHardE2E for various constraint types. The projected mean is given in Equation 8a as $\hat{\mu}=\mathcal{T}(\mu)$, and projected covariance in Equation 8b as $\hat{\Sigma}=J_{\mathcal{T}}(\mu)\Sigma J_{\mathcal{T}}(\mu)^{\top}$, where $\mathcal{T}(z)=u^{*}(z)$ denotes a transformation of the base random variable $\mathbf{Z}\sim\mathcal{F}(\mu,\Sigma)$ and $z\sim\mathbf{Z}$, and $u^{*}(z)$ denotes the solution to the constrained least squares Problem 7. For linear equality constraints, see the oblique projection $P_{Q^{-1}}$ in Equation 16; for nonlinear equality constraints, see the first-order optimality condition $R$ in Equation 19 and Blondel et al. [2022] (for implicit differentiation of $u^{*}$); and for convex inequality constraints, see Bertsekas [1997], Bonnans and Shapiro [2013] (for sensitivity analysis) and Sun et al. [2022], Agrawal et al. [2019], Amos and Kolter [2017], Gould et al. [2016] (for argmin differentiation).
Constraint Type Solution $u^{*}(z)$ Solver Type Jacobian $J_{\mathcal{T}}$
Linear Equality $P_{Q^{-1}}z+(I-P_{Q^{-1}})A^{{\dagger}}b$ closed-form $P_{Q^{-1}}$
Nonlinear Equality $(u^{*},\lambda^{*})$ s.t. $R(u^{*},\lambda^{*};z)=0\}$ nonlinear implicit differentiation
sensitivity analysis;
argmin differentiation
@@ C.1 Linear Equality Constraints
In this subsection, we provide the closed-form expressions for the constrained posterior distribution parameters, i.e., the mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ in Equation 8, from the DPPL in ProbHardE2E for linear equality constraints.
See 3.1
@@
Since the expression for $u^{*}$ in Equation 17 is a linear transformation $\mathcal{T}$ of $z\sim\mathcal{F}(\mu,\Sigma)$, we can use Theorem 3.1 with $\mathcal{T}(z)=u^{*}(z)$ to write the expression for $u^{*}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$, where:
@@ C.2 Nonlinear Equality Constraints
In this subsection, we describe how to compute the DPPL in ProbHardE2E for general nonlinear equality constraints.
See 3.2
@@
We solve Equation 19 via root-finding methods, e.g., Newton’s method for $(u^{*},\lambda^{*})$ to obtain $u^{*}(z)\;=\;\arg\{\,\hat{u}:R(\hat{u},\lambda^{*};z)=0\}\,$, where the root-finding solution $u^{*}$ is implicitly dependent on $z$. Since the expression for $u^{*}$ is a nonlinear transformation $\mathcal{T}$ of $z\sim\mathcal{F}(\mu,\Sigma)$, we can use Theorem 3.1 with $\mathcal{T}(z)=u^{*}(z)$ to write the expression for $u^{*}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$, where:
@@
$u^{*}(\mu)=\operatorname*{argmin}_{\begin{subarray}{c}\hat{u}\in\mathbb{R}^{n}\\ h(\hat{u})=0\end{subarray}}\ f(\hat{u}),$ (21)
@@
C.4 Obtaining Constrained Predictive Distribution at Inference
In our framework, we use the projected distribution parameters $(\hat{\mu},\hat{\Sigma})$ (Equation 8) obtained from the DPPL as a first-order approximation to the true constrained distribution, as justified by Theorem 3.1. The constrained mean $\hat{\mu}$ always satisfies the constraints by construction. For linear equality constraints, this affine transformation yields the exact solution (see Proposition 3.1), and the resulting distribution is fully constraint-respecting. For the more challenging nonlinear equality or convex inequality constraints, the projection map $\mathcal{T}(z)=u^{*}(z)$ is nonlinear, and the transformed distribution ${\mathcal{F}}(\hat{\mu},\hat{\Sigma})$ from Equation 8 is only first-order accurate. This sampling-free proxy suffices during training, and it enables efficient CRPS-based optimization. Due to this first-order approximation, it does not guarantee that samples drawn from $\mathcal{F}(\hat{\mu},\hat{\Sigma})$ will satisfy the nonlinear equality or convex inequality constraints exactly. A possible approach is to output the random variable $\mathbf{Z}_{\theta}(\phi^{(i)})$ in Equation 5 from the base (unconstrained) distribution, and input it to the DPPL in Equation 6 to obtain the random variable from the constrained distribution $\mathbf{Y}_{\theta}(\phi^{(i)})$. Hence, for nonlinear constraints, we impose the DPPL on the samples directly to project each sample onto the constraint manifold at inference time. This yields an empirical posterior distribution whose samples satisfy the constraints by design, even in the nonlinear case where an analytical expression for the predictive distribution is unavailable.
@@ D.1 Enforcing Coherency in Hierarchical Time Series Forecasting
[…]
@@
The projected mean and covariance for HierE2E in Equation 45 is given by the solution to Problem 7 with linear constraints in Proposition 3.1, i.e., $h(u)=Au=0$, $b=0$, where $Q=I$ for an orthogonal projection and $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ is a multivariate Gaussian.
@@ D.2 Enforcing Conservation Laws in PDEs
$\hat{\mu}_{\text{ProbConserv}}=\operatorname*{argmin}_{\begin{subarray}{c}\tilde{\mu}\in\mathbb{R}^{n}\\ A\tilde{\mu}=b\end{subarray}}\frac{1}{2}||\tilde{\mu}-\mu||^{2}_{\Sigma^{-1}}.$
We formulate this optimization problem more generally, and show that by assuming that $z\sim\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ is a multivariate Gaussian, a constrained sample $u^{*}(z)\sim\mathbf{Y}\sim\mathcal{N}(\hat{\mu},\hat{\Sigma})$ in ProbConserv is a solution to our Problem 7 with $Q=\Sigma^{-1}$ and linear constraints. In particular, we show in Proposition D.2 that the ProbConserv posterior update in Equation 48 is a special linear constraint case of our ProbHardE2E method, which uses an oblique projection with $Q=\Sigma^{-1}$.
@@
The projected mean and covariance for ProbConserv in Equation 48 is given by the solution to Problem 7 with linear constraints in Proposition 3.1, i.e., $h(u)=Au-b=0$, where $Q=\Sigma^{-1}$ for an oblique projection and $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ is a multivariate Gaussian.
@@ Appendix E Flexibility in the choice of $Q$ and its structure
where $A$ denotes a scale transformation matrix, $B$ denotes the location parameter, and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution.
@@
Then, since $\mathbf{Z}$ belongs to the location-scale family of distributions, the linearization of $\mathbf{Y}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$ also belongs to the family with mean $\hat{\mu}$ and covariance $\hat{\Sigma}$, which we compute below.
@@
[…]
$=\mathcal{T}(\mu)+J_{\mathcal{T}}(\mu)\underbrace{(\mathbb{E}[\mathbf{Z}]-\mu)}_{0}\text{ (since $\mu$ is not a random variable)}$)
@@
We note that this result is closely related to the Multivariate Delta Method [Casella and Berger, 2001], which shows that for a nonlinear function $\mathcal{T}$, the sample mean of $\mathcal{T}(z_{1},\dots,z_{n})$ also converges in distribution, under mild conditions. Specifically, if the sample mean of $n$ i.i.d. draws from $\mathbf{Z}$ converges to a multivariate Gaussian (by the CLT), then the same linearization argument and Slutsky’s theorem imply that the sample mean of the projected samples converges to a multivariate Gaussian, with parameters given in Equation 8. Second-order approximations (via a quadratic expansion of $\mathcal{T}$) yield higher-order corrections, and can lead to non-Gaussian outcomes (e.g., chi-squared) [Casella and Berger, 2001].
@@ G.1.1 Generalized Porous Medium Equation (GPME)
$k(u)$ =$=\begin{cases}1,&u\geq u^{*}\\ 0,&u<u^{*}\end{cases},\quad u^{*}\geq 0,$
@@
[…]
@@
[…]
@@ H.3 Metrics
We evaluate ProbHardE2E and the various baselines using the following metrics. We denote the exact solution or ground truth observations as $u$, and we report the metrics on the mean $\hat{\mu}$, covariance $\hat{\Sigma}$, and samples $u^{*}$ drawn from the multivariate Gaussian distribution $\mathcal{N}(\hat{\mu},\hat{\Sigma})$.
@@
$\text{MSE}(\hat{u})=\frac{1}{n}\|u-\hat{u}\|^{2}_{F},$
where the Frobenius norm is taken over all the datapoints $n$ in $u$. We report $\text{MSE}(\mu)$ on the output mean from each model.
@@
The CE measures the error in the various equality constraints $h(u)=0$, i.e., conservation laws for PDEs and coherency for hierarchical time series forecasting, and is given as:
$\text{CE}(\hat{u})=\|h(\hat{u})\|_{2}^{2}.$
For the PDEs cases with both linear and nonlinear constraints, we report the CE error on the mean $\hat{\mu}$ from each model, i.e., $\text{CE}(\hat{\mu})$. To be consistent with the evaluation results in HierE2E [Rangapuram et al., 2021], in the hierarchical forecasting case, we report $\text{CE}(u^{*})$, where $u^{*}$ denotes a sample drawn from the output distribution. Note that for linear equality constraints, $\text{CE}(\hat{\mu})=\text{CE}(u^{*})=0$ for ProbHardE2E. For more challenging nonlinear equality or convex inequality constraints, $\text{CE}(u^{*})$ is not guaranteed to be 0, due to the first order approximation with the Delta Method in Theorem 3.1. (See Section C.4 for an approach to apply the projection on the samples directly at inference time only for exact constraint enforcement.)
@@
The CRPS [Gneiting and Raftery, 2007] measures the quality of uncertainty quantification by comparing a predictive distribution to a ground-truth observation. For a multivariate Gaussian distribution with independent components $\mathcal{N}(\mu,\mathrm{diag}(\hat{\sigma}^{2}))$, where ${\hat{\sigma}}_{ii}^{2}$ denotes the $i$-th diagonal entry of the predictive covariance ${\hat{\Sigma}}$, the CRPS is given in closed-form as:
@@
[…]
@@ I.1.1 PDEs with Conservation Law Constraints
[…]
@@ I.2 Nonlinear Equality Constraints
[…]
@@
[…]
@@
where we discretize the temporal domain $[0,T]=[t_{1},\dots,t_{N_{t}}]$ into $N_{t}$ gridpoints, and TV denotes the discretized form defined in Equation 62.

### v2

@@ 5 Conclusion
In this work, we propose ProbHardE2E, which seamlessly incorporates constraints ranging from commonly used linear constraints to challenging nonlinear constraints into black-box probabilistic deep learning models using a differentiable probabilistic projection layer (DPPL). We show that ProbHardE2E is applicable in a wide range of scientific and operational domains, ranging from linear coherency constraints in time series forecasting to nonlinear conservation laws in solving PDEs. Contrary to prior works (Hansen et al., 2023), which only support linear global conservation constraints, ProbHardE2E can support nonlinear conservation constraints. This paves the way for future work to enforce conservation locally over sub-regions or over control volumes à la finite volume methods (LeVeque, 2002). Future work also includes extending our method to handle general non-convex, nonlinear inequality constraints using advanced optimization techniques or relaxations, to richer covariance parameterizations, e.g., low-rank or dense, and to empirical distributions other than location-scale families by sample projection.
@@ C.1 Linear Equality Constraints
In this subsection, we provide the closed-form expressions for the constrained posterior distribution parameters, i.e., the mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ in Equation 8, from the DPPL in ProbHardE2E for linear equality constraints. Linear equality constraints occur in a wide range of applications, including coherency constraints in hierarchical time series forecasting [Hyndman et al., 2011, Rangapuram et al., 2021, Petropoulos et al., 2022, Olivares et al., 2024b], divergence-free conditions in incompressible fluid flows [Raissi et al., 2019, Richter-Powell et al., 2022], boundary conditions in PDEs [Saad et al., 2023], and global linear conservation law constraints [Hansen et al., 2023, Mouli et al., 2024].
Proposition C.1.
[…]
@@
[…]
@@ C.2 Nonlinear Equality Constraints
In this subsection, we describe how to compute the DPPL in ProbHardE2E for general nonlinear equality constraints. Nonlinear equality constraints naturally arise in applications that involve structural, physical, or geometric consistency. These include closed-loop kinematics in robotics [Toussaint et al., 2019], nonlinear conservation laws [LeVeque, 1990] in PDE-constrained surrogate modeling [Biegler et al., 2003, Zahr and Persson, 2016, Négiar et al., 2023] with applications in climate modeling [Bolton and Zanna, 2019, Zanna and Bolton, 2020, Beucler et al., 2021], compressible flows in aerodynamics [Tezaur et al., 2017] and atomic modeling [Müller, 2022, Sturm and Wexler, 2022].
Proposition C.2.
[…]
@@
[…]
@@
[…]
@@
In this subsection, we describe how to compute the DPPL in ProbHardE2E for nonlinear convex inequality constraints. Convex inequality constraints arise naturally in many scientific and engineering applications. For example, total variation (TV) regularization is widely used to promote smoothness or piecewise-constant structure in spatial fields, e.g., image denoising [Rudin et al., 1992, Boyd and Vandenberghe, 2004] and total variation diminishing (TVD) constraints to avoid spurious artificial oscillations in numerical solutions to PDEs [Harten, 1997, Tezaur et al., 2017, Schein et al., 2021]. Other common convex constraints include box constraints, which enforce boundedness of physical or operational quantities [Bertsekas, 1997].
@@ D.1 Enforcing Coherency in Hierarchical Time Series Forecasting
Total $a_{t_{1}}$Region 1 $a_{t_{2}}$City 1 $b_{t_{1}}$City 2 $b_{t_{2}}$City 3 $b_{t_{3}}$Region 2 $a_{t_{3}}$City 4 $b_{t_{4}}$City 5 $b_{t_{5}}$City 6 $b_{t_{6}}$Region 3 $a_{t_{4}}$City 2 $b_{t_{2}}$City 3 $b_{t_{3}}$
@@
The projected mean and covariance for HierE2E in Equation 45 is given by the solution to Problem 7 with linear constraints in Proposition C.1, i.e., $h(u)=Au=0$, $b=0$, where $Q=I$ for an orthogonal projection and $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ is a multivariate Gaussian.
@@ D.2 Enforcing Conservation Laws in PDEs
[…]
@@
The projected mean and covariance for ProbConserv in Equation 48 is given by the solution to Problem 7 with linear constraints in Proposition C.1, i.e., h(u)=Aub=0h(u)=Au-b=0, where $Q=\Sigma^{-1}$ for an oblique projection and $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ is a multivariate Gaussian.
@@ Appendix E Flexibility in the choice of $Q$ and its structure
where $A$ denotes a scale transformation matrix, $B$ denotes the location parameter, and $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution.
@@
[…]
@@
$=\mathbb{E}[\mathcal{T}(\mu)]+\mathbb{E}[J_{\mathcal{T}}(\mu)(\mathbf{Z}-\mu)]\text{ (by linearity of expectation)}$
$=\mathcal{T}(\mu)+J_{\mathcal{T}}(\mu)\underbrace{(\mathbb{E}[\mathbf{Z}]-\mu)}_{0}\text{ (since $\mu$ is not a random variable)}$
@@
Importantly, the approximation error between the nonlinear transformation and its linearization converges to zero in probability [Van der Vaart, 2000], which ensures the validity of this approach asymptotically. We note that this result is closely related to the Multivariate Delta Method [Casella and Berger, 2001], which shows that for a nonlinear function $\mathcal{T}$, the sample mean of $\mathcal{T}(z_{1},\dots,z_{n})$ also converges in distribution, under mild conditions. Specifically, if the sample mean of $n$ i.i.d. draws from $\mathbf{Z}$ converges to a multivariate Gaussian (by the CLT), then the same linearization argument and Slutsky’s theorem imply that the sample mean of the projected samples converges to a multivariate Gaussian, with parameters given in Equation 8. Second-order approximations (via a quadratic expansion of $\mathcal{T}$) yield higher-order corrections, and can lead to non-Gaussian outcomes (e.g., chi-squared) [Casella and Berger, 2001].
@@ G.1.1 Generalized Porous Medium Equation (GPME)
$k(u)$ $=\begin{cases}1,&u\geq u^{*}\\ 0,&u<u^{*}\end{cases},\quad u^{*}\geq 0,$
@@
where $\mathbf{1}$ denotes the indicator function, $\mathrm{erf}(z)=(2/\sqrt{\pi})\int_{0}^{z}\exp(-y^{2})dy$ denotes the error function, and $\alpha=2\tilde{\alpha}$ and $\tilde{\alpha}$ satisfies the following nonlinear equation:
@@
$\int_{0}^{1}u(t,x)\,dx=\frac{2(1-u^{*})}{\mathrm{erf}(\alpha/2)}\sqrt{\frac{t}{\pi}},\quad\forall t\in[0,1].$ (59)
@@ H.3 Metrics
[…]
@@
$\text{MSE}(\hat{\mu})=\frac{1}{n}\|u-\hat{\mu}\|^{2}_{F},$
where the Frobenius norm is taken over all the datapoints $n$ in $\hat{\mu}$.
@@
The CE measures the error in the various equality constraints $h(u^{*})=0$, i.e., conservation laws for PDEs and coherency for hierarchical time series forecasting, on the samples, and is given as:
$\text{CE}(u^{*})=\sum_{i=1}^{N}\|h(u^{*}_{i})\|_{2}^{2},$
where we compute the average error over $N=100$ samples $\{u^{*}_{i}\}_{i=1}^{N}$.
@@
The CRPS [Gneiting and Raftery, 2007] measures the quality of uncertainty quantification by comparing a predictive distribution to a ground-truth observation. For a multivariate Gaussian distribution with independent components $\mathcal{N}(\mu,\mathrm{diag}(\hat{\sigma}^{2}))$, where ${\hat{\sigma}}_{ii}^{2}$ denotes the $i$-th diagonal entry of the predictive covariance $^{\hat{\Sigma}}$, the CRPS is given in closed-form as:
@@
where $z_{i}=(u_{i}-{\hat{\mu}}_{i})/{\hat{\sigma}}_{ii}$, $p(z_{i})=(1/\sqrt{2\pi})\exp(-z_{i}^{2}/2)$ denotes the standard normal PDF, and $P(z_{i})=\int^{z_{i}}_{-\infty}p(y)dy$ denotes the standard normal CDF [Gneiting et al., 2005, Taillardat et al., 2016].
@@ I.1.1 PDEs with Conservation Law Constraints
Figure 6 shows the analogous training time per epoch to LABEL:fig:pde_timing for PDE datasets. Models trained with 100 posterior samples per training step incur a $3.5\text{--}3.6\times$ increase in epoch time relative to our ProbHardE2E, which avoids sampling altogether by using a closed-form CRPS loss. See Table 2 for the accuracy results.
@@ I.2 Nonlinear Equality Constraints
$\int_{\Omega}u(t,x)d\Omega=\int_{0}^{t}[u^{m}(t,x_{0})\nabla u(t,x_{0})-u^{m}(t,x_{N})\nabla u(t,x_{N})]dt,\quad\forall t\in[0,1].$
@@
$\text{TV}(u(t,\cdot))=\int_{\Omega}\bigg|\frac{\partial u}{\partial x}\bigg|\,d\Omega.$
@@
where we discretize the temporal domain $0,T]=[t_{1},\dots,t_{N_{t}}]$ into $N_{t}$ gridpoints, and TV denotes the discretized form defined in Equation 62.